# Understanding Before Reasoning: Enhancing Chain-of-Thought with Iterative Summarization Pre-Prompting

### v1

@@ Abstract
Chain-of-Thought (CoT) Prompting is the dominant paradigm applied in Large Language Models (LLMs) to enhance their capacity for complex reasoning. It guides LLMs to demonstrate the problem-solving process through a chain of reasoning steps, rather than requiring LLMs to generate the final answer directly. Despite its success, CoT encounters difficulties when key information required for the reasoning process is either implicit or missing. It primarily stems from the fact that CoT emphasizes the stages of reasoning, while neglecting the critical task of gathering and extracting essential core information in the early stage. In this paper, we propose a pre-prompting methodology called Iterative Summarization Pre-Prompting ($\text{ISP}^{2}$), which can effectively refine the reasoning ability of LLMs when key information is not explicitly presented. First, entities and their corresponding descriptions are extracted to form potential key information pairs from the question. Next, we introduce the reliability rating to assess the reliability of these information pairs. Then, two information pairs with the lowest rankings through the reliability rating are merged into a new potential information description, which includes a new entity and its corresponding description. This process is applied iteratively to guide the generation of a unique information pair. Finally, the obtained key information pair, along with the original question, is fed into LLMs for reasoning, resulting in the final answer. Extensive experiments are conducted to validate the effectiveness of the proposed method. The results show that, compared to existing methods, our approach yields a 7.1% improvement in performance. In summary, unlike traditional prompting methods, $\text{ISP}^{2}$ adopts an inductive approach with pre-prompting. It demonstrates good plug-and-play performance and can theoretically be applied to improve performance across all reasoning frameworks. The code is available at: https://github.com/zdhgreat/ISP-2.
1 Introduction
Large Language Models (LLMs) have made significant strides in Natural Language Processing (NLP) tasks, such as question answering, automatic summarization, and machine translation  (Achiam et al., 2023; Chowdhery et al., 2024; Touvron et al., 2023a, b; Huang et al., 2023; Zhao et al., 2023). However, despite these advancements, LLMs still fall short in reasoning performance compared to humans. Increasing model parameters cannot close the gap in reasoning capabilities between LLMs and human intelligence. Chain-of-Thoughts (CoT) Prompting is a crucial technique for enhancing the reasoning capabilities of LLMs. It guides LLMs to break down complex reasoning into simpler steps, a process similar to human reasoning. CoT uses prompts such as "let’s think step by step" Kojima et al. (2022) and multiple learnable examples Wei et al. (2022) to generate interpretable prediction paths. It has made LLMs excellent zero-shot and few-shot reasoners, paving the way for tackling complex problems. When faced with more complex problems, CoT prompting methods often fail to address them effectively. During the reasoning process, some information is often overlooked, which results in unclear guiding strategies and ultimately impacts the quality of the answer.
@@ 1 Introduction
Simon (1978)’s theory provides valuable insights to better understand and solve problems. First, individuals gather key information from the context of the problem and use it to construct the problem space. Then, with an understanding of the current information, they use heuristic strategies to summarize and refine their thoughts, gradually approaching the solution to the problem. Building upon Simon’s foundational work in information processing theory, this paper applies these concepts to the reasoning of LLMs and introduces Iterative Summarization Prompting ($\text{ISP}^{2}$). It leverages the inherent knowledge of LLMs to develop precise descriptions of specific problem spaces. Specifically, $\text{ISP}^{2}$ is a pre-prompting method applied before CoT, allowing the LLMs to summarize more comprehensive knowledge to assist in reasoning. In Figure 1, we illustrate the workflow differences between standard prompting and the pre-prompting method. It coordinates three key LLM steps: adaptive extraction of candidate information, reliability rating of information pairs, and iterative summarization for knowledge understanding. These steps include summarizing and integrating relevant information, as well as formulating strategies before tackling intricate real-world reasoning tasks. By engaging in these pre-prompting steps, LLMs can better explore and understand the nuances of complex problems, thereby improving their ability to perform sophisticated reasoning. Testing with GPT-3.5 Turbo shows that inserting $\text{ISP}^{2}$ before CoT and Complex CoT leads to a significant performance improvement, with increases of 7.1% and 8.1%, respectively. The average performance score of $\text{ISP}^{2}$ with CoT reaches 79.43, surpassing other SOTA methods with plug-and-play capabilities. In these processes and results, our main contributions are as follows:
Figure 1: The difference between prompting and pre-prompting. Conventional prompting focuses on providing reinforcement during the inference process to guide the model’s reasoning. In contrast, pre-prompting enriches the input context prior to reasoning, ensuring that the model has a more comprehensive and refined understanding from the outset.
@@
  • We introduce a pre-prompting method that can be seamlessly integrated into various CoT methods to enhance their reasoning performance.
@@
  • $\text{ISP}^{2}$ demonstrates robustness, performing exceptionally well across diverse models, including GPT-3.5 Turbo and open-source LLaMA2, proving its effectiveness in various reasoning environments. Furthermore, it achieves top performance among SOTA plug-and-play methods.
2 Related Work
2.1 Chain-of-Thoughts Prompting
@@
Wei et al. (2022) emphasize the importance of deriving conclusive answers through multi-step logical pathways by introducing the concept of Chain-of-Thoughts (CoT) reasoning. The method demonstrates that reasoning abilities can be elicited through a series of thoughtful steps. Kojima et al. (2022) discover that simply adding the phrase "let’s think step by step" in prompts allows LLMs to perform zero-shot logical reasoning without any additional human prompts. Subsequently, Wang et al. (2023) introduce Self Consistency (SC) to replace the greedy decoding strategy. Zhang et al. (2023) construct an automatic CoT framework based on the problem, eliminating the instability of manual prompts. Fu et al. (2023) employ complexity-based multi-step reasoning estimation to execute CoT. Yao et al. (2024) propose Tree-of-Thoughts (ToT), which introduces deliberation into decision-making by considering multiple reasoning paths. Xu et al. (2024) enhance the model’s understanding by re-reading the question. These studies underscore the importance of CoT in enhancing the reasoning and planning capabilities of LLMs in complex scenarios. Despite, CoT still requires further refinement in complex scenarios involving more complex problems.
2.2 In-Context Learning
@@
In-context learning (ICL) enables LLMs to make predictions based on input examples without updating model parameters. Brown et al. (2020) introduce this concept in GPT-3, demonstrating that LLMs can generalize tasks from a small number of examples embedded in the input context. Min et al. (2022) propose Meta-training for In-Context Learning (MetaICL), which significantly enhances ICL capabilities through continuous training on various tasks using demonstrations. Additionally, the concept of supervised context trainingChen et al. (2022) is proposed to bridge the gap between pre-training and downstream ICL tasks. LLM refines its prior knowledge through ICL, thereby improving performance across multiple tasks Krishnamurthy et al. (2024). ICL allows a single model to perform various tasks universally, helping it better align its predictions with the semantic requirements of the prompts.
2.3 Task Decomposition
@@
Perez et al. (2020) decompose complex problems into several independent subproblems by the LLMs, and then aggregates the answers to form the final response. Wang, Deng, and Sun (2022) address problems by modeling prompts as continuous virtual tokens and iteratively eliciting relevant knowledge from a LLM. Yang et al. (2022) decompose normal questions into a series of subproblems, which are then converted into SQL queries using a rule-based system. Wu, Terry, and Cai (2022) introduce the idea of linking LLM steps, where the output of one step becomes the input of the next, and developed an interactive system for users to build and modify these chains. Zhou et al. (2023) argue that generated subproblems are often interdependent and need to be solved in a specific order, with the answers to some subproblems serving as the foundation for others. They propose the Least-to-Most Prompting method, which links the problem decomposition process to the solving of subproblems. Zhang et al. (2024) propose the Cumulative Reasoning (CR), breaking down complex tasks into smaller manageable steps and utilizing iterative collaboration among three different LLMs to incrementally solve problems.
2.4 Self Evaluation
@@
Researchers have proposed automated evaluation methods, such as Sentence-BERT Reimers and Gurevych (2019) and SimCSE Gao, Yao, and Chen (2021), to assess the reasoning process. However, these methods primarily concentrate on matching individual words and phrases, which limits their ability to fully assess the logical consistency and deeper meaning of the context. To address these limitations, the feasibility of using LLMs to evaluate their own predictions is becoming an increasingly important step in problem-solving. Shinn et al. (2023), Madaan et al. (2024) and Paul et al. (2024) introduce the Self Evaluation (SE) mechanism, where LLMs provide feedback on the candidate answers they generate. Chen et al. (2024) improves LLM code generation accuracy by using self-generated feedback. Similarly, Kim, Baldi, and McAleer (2024) introduce a review step to evaluate actions and states in operational tasks and decide the next steps. In terms of reasoning, Yao et al. (2024) emphasize SE guided decoding, where the LLM uses carefully designed prompts to evaluate candidate answers via a tree search procedure. Kumar et al. (2024) explore facilitate scalable self-reflection in LLM, demonstrating its effectiveness in improving student learning outcomes. By incorporating fair assessment in LLM learning, our approach injects the reflection mechanism into problem space understanding rather than just evaluating candidate answers, allowing for deeper consideration of the problem and focusing more on the essence of the problem. We believe that reasoning based on a thorough understanding of information leads to further refinement and improvement.
3 Proposed Approach
@@
The key to problem-solving lies in thoroughly understanding the problem and integrating important information based on the context. The process of understanding involves not only identifying the key parts that contain the answers but also clearly representing the problem space. The deep understanding helps reduce cognitive load and accelerates the process of finding answers. In this work, we enhance reasoning by constructing knowledge representations of the problem space, grounded in a thorough understanding of the question at hand. As illustrated in Figure 2, we propose Iterative Summarization Pre-Prompting ($\text{ISP}^{2}$), a plug-and-play pre-prompting method that operates in three steps: adaptive extraction of candidate information, reliability rating of information pairs, and iterative summarization for knowledge understanding. Throughout the progressive process, the LLM continuously enhances its understanding of the information, thereby more effectively exploring and solving problems.
3.1 Adaptive Extraction of Candidate Information
@@
LLMs demonstrate distinct advantages over traditional machine learning methods in extracting and synthesizing complex information, thanks to their sophisticated architectures and extensive training on broad datasets Schacht, Kamath Barkur, and Lanquillon (2023). Given the strong information extraction capabilities, we leverage the LLM to extract multiple entities $E$ under the question $Q$. To effectively understand a problem, it is crucial to gather all relevant information related to the question as candidates. The comprehensive collection of potential information forms the foundation for subsequent deepening of understanding. During the extraction process, not all entities maintain a high degree of relevance to $Q$. Therefore, we design prompts and provide sufficient examples to guide the LLM. It allows the LLM to learn from the examples in the prompt $T$ and effectively retain the $n$ entities $E=\{(e_{i})\}_{i=1}^{n}$ that are highly relevant to the question. For each entity, the LLM summarizes the knowledge descriptions $K=\{k_{i1},k_{i2},\ldots,k_{it}\}_{i=1}^{n}$ that are related to it. The entities and their corresponding knowledge descriptions are organized into information pairs $IP=\left[e_{i},\{k_{i1},k_{i2},\ldots,k_{it}\}\right]_{i=1}^{n}$. Each information pair focuses on summarizing explicit knowledge in the question, laying the groundwork for deeper analysis. Meanwhile, we do not specify the number of entities or knowledge descriptions due to the varying amount of extracted information for each question. This flexible approach ensures that the LLM can adapt to the specific requirements of different problems, continuously enhancing its understanding and refining its outputs.
3.2 Reliability Rating of Information Pairs
Inspired by the recent success of Self Evaluation Kadavath et al. (2022), we introduce an automated evaluation method called reliability rating. A key aspect of understanding lies in uncovering potentially lost information, and reliability rating plays a crucial role in this process. Reliability rating not only assesses the potential of each information pair to contribute effectively to problem-solving but also evaluates the completeness of the candidate information. It evaluates the current information pair $IP_{t}$ by considering all previously generated information pairs $IP_{1:n}$. We define the value of each information pair using the function $V(IP_{t})$:
$V(IP_{t})\sim p(v\mid T,Q,IP_{1:n}),$ (1)
which quantifies its contribution to solving the problem while ensuring the integrity of the information. Importantly, a lower reliability score for an information pair indicates a higher likelihood of missing critical information. The low-scored pairs are prioritized for early inclusion in the iterative summarization process, allowing for the gradual supplementation of missing details through successive iterations. By evaluating the potential and completeness of each information pair, the method refines and prioritizes the information used in the summarization process, thereby enhancing the overall effectiveness and accuracy of the solution. where the prompts $T$ for evaluating information pairs are divided into two types:

### v2

@@ Abstract
Chain-of-Thought (CoT) is the dominant paradigm applied in Large Language Models (LLMs) to enhance their capacity for complex reasoning. It guides LLMs to demonstrate the problem-solving process through a chain of reasoning steps, rather than requiring LLMs to generate the final answer directly. Despite its success, CoT encounters difficulties when key information required for the reasoning process is either implicit or missing. It primarily stems from the fact that CoT emphasizes the stages of reasoning, while neglecting the critical task of gathering and extracting essential core information in the early stage. In this paper, we propose a pre-prompting methodology called Iterative Summarization Pre-Prompting ($\text{ISP}^{2}$), which can effectively refine the reasoning ability of LLMs when key information is not explicitly presented. First, entities and their corresponding descriptions are extracted to form potential key information pairs from the question. Next, we introduce the reliability rating to assess the reliability of these information pairs, prioritizing those with lower rankings which retain effectiveness for problem-solving but require further knowledge extraction. Then, the two lowest-ranked pairs, which demonstrate problem-solving capacity yet underdeveloped knowledge associations, are merged through summarization. The summarization process synthesizes two pairs into a consolidated description, exposing implicit relationships that exist between information pairs. The reliability rating and summarization are iteratively applied to guide the generation of a unique information pair. Finally, the obtained unique information pair, along with the original question, is fed into LLMs for reasoning, resulting in the final answer. Extensive experiments are conducted to validate the effectiveness of the proposed method. The results show that, compared to existing methods, our approach yields a 7.1% improvement in performance. In summary, unlike traditional prompting methods, $\text{ISP}^{2}$ adopts an inductive approach with pre-prompting. It demonstrates good plug-and-play performance and can theoretically be applied to improve performance across all reasoning frameworks. The code is available at: https://github.com/X-Lab-CN/ISP.
I Introduction
Large Language Models (LLMs) have made significant strides in Natural Language Processing (NLP) tasks, such as question answering, automatic summarization, and machine translation [1, 2, 3, 4, 5, 6]. However, merely scaling up model parameters has shown limited effectiveness in bridging the gap between LLM reasoning capabilities and human-level performance. Chain-of-Thought (CoT) has emerged as a promising approach to enhance reasoning processes. By decomposing complex problems into intermediate steps through prompts like ”let’s think step by step” [7] or demonstration-based learning [8], CoT has demonstrated measurable improvements in zero-shot and few-shot reasoning benchmarks. Nevertheless, the effectiveness of CoT methods depends on the complexity of the problem. While CoT provides more interpretable prediction paths for moderately complex tasks, it faces limitations when handling highly complex reasoning scenarios. Challenges such as the frequent overlooking of critical information can lead to unclear guiding strategies, ultimately impacting the quality of the answers.
Simon et al.[9]’s theory provides valuable insights to better understand and solve problems. The theory establishes a formal framework for problem-solving analysis, where human cognition constructs problem spaces through systematic operator applications and goal formalization. With an understanding of the current context, humans use heuristic strategies to summarize and refine their thoughts, gradually approaching the solution to the goal. Building upon Simon’s foundational work in information processing theory, this paper applies these concepts to the reasoning of LLMs and introduces Iterative Summarization Pre-Prompting ($\text{ISP}^{2}$). $\text{ISP}^{2}$ operationalizes three core processes aligned with classical problem-solving theory: (1) formalization of state transition operators via task-driven input filtering, (2) rule-based evaluation of relational mappings under dynamic constraints, and (3) goal-bound hierarchical consolidation with termination conditions. The structured method addresses the limitations of implicit knowledge assumptions in LLMs by explicitly constructing problem representations through observable operations, thereby maintaining theoretical consistency with Simon’s information processing work.
@@ I Introduction
Specifically, $\text{ISP}^{2}$ is a pre-prompting method applied before CoT, allowing the LLMs to summarize more comprehensive knowledge to assist in reasoning. In Figure 1, we illustrate the differences between two paradigms for enhancing CoT reasoning: one is the mainstream approach that augments CoT during the reasoning stream, and the other is our method of enhancing CoT through pre-prompting. $\text{ISP}^{2}$ enhances CoT through pre-prompting by systematically enriching the input context before reasoning. It coordinates three key LLM steps: adaptive extraction of candidate information, reliability rating of information pairs, and iterative summarization for knowledge understanding. These steps include summarizing and integrating relevant information, as well as formulating strategies before tackling intricate real-world reasoning tasks. By engaging in these pre-prompting steps, LLMs can better explore and understand the nuances of complex problems, thereby improving their ability to perform sophisticated reasoning. Testing with GPT-3.5 Turbo shows that inserting $\text{ISP}^{2}$ leads to a significant performance improvement, with increases of 7.1% and 8.1%, respectively. The average performance score of $\text{ISP}^{2}$ with CoT reaches 79.43, surpassing other SOTA methods with plug-and-play capabilities. The results validate the effectiveness of integrating formal problem space principles into modern LLM prompting strategies. In these processes and results, our main contributions are as follows:
Figure 1: The difference between mainstream prompt enhanced CoT and pre-prompt enhanced CoT. The left part illustrates the mainstream approach, which augments CoT during the reasoning process by dynamically guiding the model’s inference steps. The right part demonstrates the pre-prompt method, which enhances CoT through pre-prompting by systematically enriching the input context before reasoning. The distinction highlights how pre-prompting shifts the focus from in-process guidance to upfront contextual refinement, enabling more robust reasoning from the outset.
@@
  • $\text{ISP}^{2}$ is a pre-prompting method that can be seamlessly integrated into various CoT methods to enhance their reasoning performance. Furthermore, it achieves top performance among SOTA plug-and-play methods.
@@
  • The results demonstrate that $\text{ISP}^{2}$ achieves outstanding performance across various LLM environments, from closed-source models (GPT-3.5 Turbo, GPT-4o-mini) to open-source models (LLaMA2/3, DeepSeek-Distill), verifying its compatibility with diverse parameter regimes and transformer-based architectures.
II Related Work
II-A Chain-of-Thoughts Prompting
@@
Wei et al.[8] emphasize the importance of deriving conclusive answers through multi-step logical pathways by introducing the concept of Chain-of-Thoughts (CoT) reasoning. The method demonstrates that reasoning abilities can be elicited through a series of thoughtful steps. Kojima et al.[7] discover that simply adding the phrase ”let’s think step by step” in prompts allows LLMs to perform zero-shot logical reasoning without any additional human prompts. Subsequently, Wang et al.[10] introduce Self Consistency (SC) to replace the greedy decoding strategy. Zhang et al.[11] construct an automatic CoT framework based on the problem, eliminating the instability of manual prompts. Fu et al.[12] employ complexity-based multi-step reasoning estimation to execute CoT. Yao et al.[13] propose Tree-of-Thoughts (ToT), which introduces deliberation into decision-making by considering multiple reasoning paths. Xu et al.[14] enhance the model’s understanding by re-reading the question. These studies underscore the importance of CoT in enhancing the reasoning and planning capabilities of LLMs in complex scenarios. Despite, CoT still requires further refinement in complex scenarios involving more complex problems.
II-B In-Context Learning
@@
In-context learning (ICL) enables LLMs to make predictions based on input examples without updating model parameters. Brown et al.[15] introduce this concept in GPT-3, demonstrating that LLMs can generalize tasks from a small number of examples embedded in the input context. Min et al.[16] propose Meta-training for In-Context Learning (MetaICL), which significantly enhances ICL capabilities through continuous training on various tasks using demonstrations. Additionally, the concept of supervised context training[17] is proposed to bridge the gap between pre-training and downstream ICL tasks. LLM refines its prior knowledge through ICL, thereby improving performance across multiple tasks [18]. ICL allows a single model to perform various tasks universally, helping it better align its predictions with the semantic requirements of the prompts.
II-C Task Decomposition
@@
Perez et al.[19] decompose complex problems into several independent subproblems by the LLMs, and then aggregates the answers to form the final response. Wang et al.[20] address problems by modeling prompts as continuous virtual tokens and iteratively eliciting relevant knowledge from a LLM. Yang et al.[21] decompose normal questions into a series of subproblems, which are then converted into SQL queries using a rule-based system. Wu et al.[22] introduce the idea of linking LLM steps, where the output of one step becomes the input of the next, and developed an interactive system for users to build and modify these chains. Zhou et al.[23] argue that generated subproblems are often interdependent and need to be solved in a specific order, with the answers to some subproblems serving as the foundation for others. They propose the Least-to-Most Prompting method, which links the problem decomposition process to the solving of subproblems. Zhang et al.[24] propose the Cumulative Reasoning (CR), breaking down complex tasks into smaller manageable steps and utilizing iterative collaboration among three different LLMs to incrementally solve problems.
II-D Self Evaluation
@@
Researchers have proposed automated evaluation methods, such as Sentence-BERT [25] and SimCSE [26], to assess the reasoning process. However, these methods primarily concentrate on matching individual words and phrases, which limits their ability to fully assess the logical consistency and deeper meaning of the context. To address these limitations, the feasibility of using LLMs to evaluate their own predictions is becoming an increasingly important step in problem-solving. Shinn et al.[27], Kim et al.[28] and Paul et al.[29] introduce the Self Evaluation (SE) mechanism, where LLMs provide feedback on the candidate answers they generate. Chen et al.[30] improve LLM code generation accuracy by using self-generated feedback. Similarly, Kim et al.[31] introduce a review step to evaluate actions and states in operational tasks and decide the next steps. In terms of reasoning, Kim et al.[13] emphasize SE guided decoding, where the LLM uses carefully designed prompts to evaluate candidate answers via a tree search procedure. Kumar et al.[32] explore facilitate scalable self-reflection in LLM, demonstrating its effectiveness in improving student learning outcomes. By incorporating fair assessment in LLM learning, our approach injects the reflection mechanism into problem space understanding rather than just evaluating candidate answers, allowing for deeper consideration of the problem and focusing more on the essence of the problem.
III Proposed Approach
@@
The key to problem-solving depends on rigorous semantic parsing of task requirements and systematic integration of contextual information. The process requires both precise identification of the key components that contain the answers and formal representation of the problem space through structured knowledge. Such formalization reduces the cognitive complexity of problem-solving while explicitly constructing the problem space. As illustrated in Figure 2, we propose Iterative Summarization Pre-Prompting ($\text{ISP}^{2}$), a plug-and-play pre-prompting method that operates in three steps: adaptive extraction of candidate information, reliability rating of information pairs, and iterative summarization for knowledge understanding. $\text{ISP}^{2}$ enables continuous optimization of knowledge representations through successive approximation, thereby enhancing the model’s capacity for problem resolution.
III-A Adaptive Extraction of Candidate Information
@@
LLMs demonstrate distinct advantages over traditional machine learning methods in extracting and synthesizing complex information, thanks to their sophisticated architectures and extensive training on broad datasets [33]. Given the strong information extraction capabilities, we leverage the LLM to extract multiple entities $E$ under the question $Q$. To ensure extraction relevance, the prompt adopts example-driven learning, providing two template extraction cases in the prompt. The template instructions are accompanied by demonstrations of entity extraction for questions in the relevant domain. It enables the LLM to learn the relevance criteria for extraction across examples (such as semantic proximity and causal influence), ultimately mimicking the demonstrated behavior to extract the final entity set $E=\{e_{i}\}_{i=1}^{n}$ without requiring an explicit scoring stage. For each retained entity, the LLM also generates structured information descriptions $K=\{k_{i1},k_{i2},\ldots,k_{it}\}_{i=1}^{n}$ through a template-driven process. These descriptions typically include definitional attributes (e.g., physical laws), contextual constraints (e.g., applicable domains), and operational relationships (e.g., mathematical dependencies). The entities and their corresponding information descriptions are organized into information pairs $IP=\left[e_{i},\{k_{i1},k_{i2},\ldots,k_{it}\}\right]_{i=1}^{n}$. This step of acquiring information pairs is called adaptive extraction, where each information pair focuses on summarizing the explicit information in the problem. Adaptive extraction lays the foundation for subsequent iterative summarization, where explicit information serves as an anchor for inferring implicit information through structured knowledge propagation. Meanwhile, we do not specify the number of entities or information descriptions due to the varying amount of extracted information for each question. This flexible approach ensures that the LLM can adapt to the specific requirements of different problems, continuously enhancing its understanding and refining its outputs. The entity and description prompts are structured as follows:
Entity Prompting {2 Entity Template}
Q: {Input Query}
Generate key entities: # Thought-eliciting prompt (e.g.,‘‘Let’s think step by step") #
Description Prompting {2 Description Template}
Q: {Input Query and Entities}
Generate descriptions: # Thought-eliciting prompt (e.g.,‘‘Let’s think step by step") #
III-B Reliability Rating of Information Pairs
Inspired by the recent success of Self Evaluation [34], we introduce an automated evaluation method called reliability rating. It quantifies the quality and reliability of information pairs, assessing their problem-solving potential and completeness. It evaluates the current information pair $IP_{t}$ conditioned on all preceding pairs $IP_{1:n}$. By leveraging a LLM, we compute the reliability rating as an expectation over the model’s probability distribution:
$V(IP_{t})=\mathbb{E}_{v\sim p_{\text{LLM}}(v\mid T,Q,IP_{1:n})}[v],$ (1)
where $p_{\text{LLM}}(v\mid\cdot)$ represents the LLM’s probabilistic assessment of $IP_{t}$’s utility. Conditioning on $T$, $Q$, and $IP_{1:n}$ allows reliability rating to integrate contextual dependencies into the evaluation process. It ensures that the reliability score reflects not merely a single deterministic outcome, but a comprehensive assessment of the information’s quality within the problem-solving context. By anchoring the evaluation in the interplay between task-specific requirements, question, and the evolving information pairs, the reliability rating becomes sensitive to the logical coherence and relevance of information pairs. The expectation-based evaluation method makes integrate diverse factors contributing to reliability(such as fluency, relevance, and completeness)into a unified score. Rather than relying on a rigid weighted sum of individual dimensions, this method leverages the LLM’s learned representations to balance these factors dynamically, resulting in a more flexible and interpretable metric. LLMs inherently capture the uncertainty and variability in natural language generation, and their probability distributions provide a principled way to quantify the likelihood of different outputs. In scenarios where exact computation of the expectation $\mathbb{E}_{v\sim p_{\text{LLM}}(v\mid T,Q,IP_{1:n})}[v]$ is infeasible or computationally expensive, we approximate it using multiple sampling. Specifically, we draw $k$ independent samples $v^{(i)}$ from the conditional probability distribution $p_{\text{LLM}}(v\mid T,Q,IP_{1:n})$ and compute the reliability rating as the empirical average:
$V(IP_{t})=\frac{1}{k}\sum_{i=1}^{k}v^{(i)},\quad v^{(i)}\sim p_{\text{LLM}}(v\mid T,Q,IP_{1:n}),$ (2)
where $v^{(i)}$ represents the score assigned to the $i$-th sample. The number of samples $k$ controls the trade-off between estimation accuracy and computational cost. Given the experimental constraints, we set the number of samples to 3.
The prompts $T$ for evaluating information pairs are divided into two types: